For the following cash flows, compute the following assuming a 10% interest rate: PW EAW FW...

  1. For the following cash flows, compute the following assuming a 10% interest rate:
    1. PW
    2. EAW
    3. FW
    4. Discounted payback period (year at which the sum of or cumulative PW of cash flows in years 1 through N = initial investment)
    5. Benefit / Cost Ratio
    6. Find Internal rate of return of the cash flows

Year

Cash Flow

0

-4000

1

250

2

500

3

750

4

1000

5

1250

6

1500
Answer #1
  1. Present Worth (PW)=Future Value/(1+i)^n

Therefore Present Worth=

1(-4000)+250(0.909)+500(0.826)+750(0.751)+1000(0.683)+1250(0.621)+1500(0.564)

=-4000+227.25+413+563.25+683+776.25+846=-$904.25

  1. EAW=rate of interest(Net present value)/1-(1+r)^-n

=$ 904.25 (0.10)/{1-(1+0.10)^-4

=$904.25(0.10)/0.31698654

=285.264

  1. FutureWorth=Present value(1+rate of interest)^n

(4000)(1)+227.25(1.10)+413(1.10)^2+563.25(1.10)^3+683(1.10)^4+776.25(1.10)^5+846(1.10)^6

=-4000+249.975+499.73+749.686+999.980+1250.158+1498.741

=$1248.27

  1. Discounted payback period= Beyond 6 years

Year

Cash Flow

Cummulative value of

Present value of

Cash flow @10%

0

(-4000)

-4000

1

250

227.25=227.25

2

500

227.25+413=640.25

3

750

640.25+563.25=1203.5

4

1000

1203.5+683=1886.5

5

1250

1886.5+776.25=2662.75

6`

1500

2662.75+846=3508.75

Discounted payback is approximately beyond 6 years